OpenStudy (anonymous):

Analyze the equation of the ellipse: 4x^2+3y^2+16x-30y=-55 I got to here and I am stuck now. 4(x^2+4x)+3(y^2-10y)=-55

OpenStudy (anonymous):

complete the square for each part

OpenStudy (anonymous):

i'm noy sure of how to do that

OpenStudy (anonymous):

\[4(x^2+4x+4)=3(y^2-10y+25)=-55+4\times 4+3\times 25\] is the first step

OpenStudy (anonymous):

arithmetic on the right the point of adding those numbers is now you can write \[4(x+2)^2+3(x-5)^2=36\]

OpenStudy (anonymous):

now divide by \(36\) and get \[\frac{(x+2)^2}{9}+\frac{(y-5)^2}{12}=1\]

OpenStudy (anonymous):

i see i made lots of typos, but the last line is correct you know how to complete the square?

OpenStudy (anonymous):

By dividing what was factored out ?

OpenStudy (anonymous):

you mean the step from here \[4(x+2)^2+3(y-5)^2=36\] to here\[\frac{(x+2)^2}{9}+\frac{(y-5)^2}{12}=1\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

divided by 36 to get a 1 on the right and reduced to lowest terms

OpenStudy (anonymous):

\[\frac{4}{36}=\frac{1}{9}\\ \frac{3}{36}=\frac{1}{12}\]

OpenStudy (anonymous):

oh ok. I also have to find the center, vertices endpoints and foci. Do I use what you got at the end ?

OpenStudy (anonymous):

yes